OpenStudy (anonymous):

Identify intervals on which the function is increasing, decreasing, or constant. g(x) = 4 - (x - 6)^2

OpenStudy (anonymous):

help meee

OpenStudy (anonymous):

@RadEn can you help me??

OpenStudy (raden):

is it calculus ?

OpenStudy (anonymous):

precalc!

OpenStudy (raden):

have u learned about differential ?

OpenStudy (anonymous):

a. Increasing: x < 4; decreasing: x > 4 b. Increasing: x < 6; decreasing: x > 6 c. Increasing: x < -6; decreasing: x > -6 d. Increasing: x > 4; decreasing: x < 4

OpenStudy (anonymous):

not suree

OpenStudy (raden):

use differential, take the first derivative of g(x) so, if g(x) = 4 - (x - 6)^2 then g ' (x) = ...

OpenStudy (raden):

hint : use the chain rule

OpenStudy (raden):

can u ?

OpenStudy (anonymous):

umm not sure how to do it...

OpenStudy (raden):

well, to get g' the other way is we simply the g(x), first g(x) = 4 - (x - 6)^2 = 4 - (x^2 - 12x + 36) = -x^2 + 12x - 32

OpenStudy (raden):

so, if g(x) = -x^2 + 12x - 32 then g' = -2x + 12, right ?

OpenStudy (anonymous):

yes!

OpenStudy (raden):

now, g(x) would increasing when g' > 0 and decreasing when g'(x) < 0

OpenStudy (raden):

case I : g'(x) > 0 -2x + 12 > 0 -2x > -12 x < 6 this, intervals when g(x) would increas

OpenStudy (raden):

case II : g'(x) < 0 -2x + 12 < 0 -2x < -12 x > 6 this, intervals when g(x) would decreas

OpenStudy (anonymous):

ohh so that makes the answer B right? Increasing: x < 6; decreasing: x > 6

OpenStudy (raden):

yes

OpenStudy (anonymous):

ahh i get it thank you!!!!!!:)

OpenStudy (raden):

you're welcome